The Magellanic Clouds Newsletter

An electronic exchange on Magellanic Clouds research

Edited by Eva K. Grebel and You-Hua Chu

                                    mcnews@astro.uiuc.edu
                         http://www.astro.uiuc.edu/mcnews/MCNews.html
                            http://www.astro.uni-bonn.de/~mcnews/

No. 30 March 7, 1999

Contents

Abstracts of 12 refereed papers  1
Abstract of a non-refereed paper 9
Job opportunity  10

Abstracts of Refereed Papers

The orientation within the Galaxy and the Large Magellanic Cloud of nebulae ejected by massive stars

D. Hutsemékers

Liège Astrophysical Institute, Belgium

The orientation of nebulae ejected by massive stars (Luminous Blue Variables, WR stars, SN1987A) is investigated with respect to the structure of the galaxy to which they belong.

In the Galaxy, we find that the projected long axes of the nebulae most often align with the galactic plane, and then also with the galactic magnetic field. This alignment is statistically significant. In addition, a few nebulae are apparently oriented perpendicular to the galactic plane. In the Large Magellanic Cloud, the nebular axes are found to closely follow the spiral magnetic field. With different inclinations, the Galaxy and the Large Magellanic Cloud probably offer complementary views of the same phenomenon.

Although the sample studied thus far is small and the statistics limited, these results suggest that the orientation of massive star ejecta depends on galactic magnetic fields.

Since the nebular axes are apparently correlated to the symmetry axes of the stars themselves, and since, in the early evolutionary stages, alignments of accretion disk axes with the interstellar magnetic field have been reported, it is argued that the observed alignment effect results from the star formation process.

Accepted by: Astronomy & Astrophysics

For preprints, contact hutsemek@astro.ulg.ac.be
Also available from the URL http://vela.astro.ulg.ac.be/preprint/P34/index.html

HST/FOS Spatially Resolved Spectral Classification of Compact OB Groups in the LMC

Nolan R. Walborn^{1}, Laurent Drissen^{2}, Joel Wm. Parker^{3}, Abhijit Saha^{4}, John W. MacKenty^{1}, and Richard L. White^{1}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2 Département de Physique, Université Laval, Ste-Foy, Québec G1K 7P4, Canada 3 Southwest Research Institute, 1050 Walnut Street, Suite 426, Boulder, CO 80302, USA 4 Kitt Peak National Observatory, National Optical Astronomy Observatories, 950 North Cherry Avenue, Tucson, AZ 85726, USA

Blue-violet spectrograms of individual components in four compact OB groups of the Large Magellanic Cloud, obtained with the Hubble Space Telescope Faint Object Spectrograph, are presented and discussed. Two of the massive multiple systems are in the 30 Doradus periphery, while the other two represent the core and the peripheral, triggered associations in the giant shell H ii region Henize N11. (More specifically, these systems are the WN + OB groups Breysacher 73 in 30 Doradus B and NGC 2044 West (HDE 269828) in 30 Doradus C; and the central WC + OB object HD 32228 in N11, along with the O3 III(f*) + OB group Lucke-Hodge 10/Parker 3209 in the largest nebula of the N11 shell.) Uncontaminated spectrograms of three WR and two very early Of components have been obtained for the first time; they can be observed only as composites with their close companions from the ground. Many of the companions have also been observed separately with HST/FOS, and several are of special interest in their own right. In particular, an early O giant in Brey 73 has twice the mass of the apparent turnoff of the compact group, and it is proposed as a possible massive stellar merger, perhaps providing a clue to the interpretation of the peculiar spectroscopic category to which it belongs. These observations provide information on the initial masses and ages of the peculiar objects, and on the evolutionary relationships among different spectral categories within the presumably coeval systems. The results are also relevant to the upper stellar mass limit and initial mass function.

Submitted to: The Astronomical Journal For preprints, contact walborn@stsci.edu

IRAS04496−6958: A luminous carbon star with silicate dust in the Large Magellanic Cloud

Norman R. Trams^{1}, Jacco Th. van Loon^{2}, Albert A. Zijlstra^{3}, Cecile Loup^{4}, M.A.T. Groenewegen^{5}, L.B.F.M. Waters^{2,6}, Patricia A. Whitelock^{7}, J.A.D.L. Blommaert^{8}, Ralf Siebenmorgen^{8} and Astrid Heske^{8}

1 Integral Science Operations Centre, Astroph. Div., Science Dep., ESTEC, P.O.Box 299, NL-2200 AG Noordwijk, The Netherlands 2 Astronomical Institute, University of Amsterdam, Kruislaan 403, NL-1098 SJ Amsterdam, The Netherlands 2 University of Manchester Institute of Science and Technology, P.O.Box 88, Manchester M60 1QD, United Kingdom 4 Institut d’Astrophysique de Paris, 98bis Boulevard Arago, F-75014 Paris, France 5 Max-Planck Institut für Astrophysik, Karl-Schwarzschild Straße 1, D-85740 Garching bei München, Germany 6 Space Research Organization Netherlands, Landleven 12, NL-9700 AV Groningen, The Netherlands 7 South African Astronomical Observatory, P.O.Box 9, 7935 Observatory, South Africa 8 ISO Data Centre, Astroph. Div., Science Dept. of ESA, Villafranca del Castillo, P.O.Box 50727, E-28080 Madrid, Spain

We describe ISO observations of the obscured Asymptotic Giant Branch (AGB) star IRAS04496−6958 in the Large Magellanic Cloud (LMC). This star has been classified as a carbon star. Our new ISO-CAM CVF spectra show that it is the first carbon star with silicate dust known outside of the Milky Way. The existence of this object, and the fact that it is one of the highest luminosity AGB stars in the LMC, provide important information for theoretical models of AGB evolution and understanding the origin of silicate carbon stars.

Accepted by: Astronomy and Astrophysics Letters

For preprints, contact ntrams@astro.estec.esa.nl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9902323

Luminous carbon stars in the Large Magellanic Cloud

Jacco Th. van Loon^{1}, Albert A. Zijlstra^{2} and M.A.T. Groenewegen^{3}

1 Astronomical Institute, University of Amsterdam, Kruislaan 403, NL-1098 SJ Amsterdam, The Netherlands 2 University of Manchester Institute of Science and Technology, P.O.Box 88, Manchester M60 1QD, United Kingdom 3 Max-Planck Institut für Astrophysik, Karl-Schwarzschild Straße 1, D-85740 Garching bei München, Germany

We present ground-based 3 μm spectra of obscured Asymptotic Giant Branch (AGB) stars in the Magellanic Clouds (MCs). We identify the carbon stars on the basis of the 3.1 μm absorption by HCN and C_{2}H_{2} molecules.

We show evidence for the existence of carbon stars up to the highest AGB luminosities (M_{bol} = −7 mag, for a distance modulus to the LMC of 18.7 mag). This proves that Hot Bottom Burning (HBB) cannot, in itself, prevent massive AGB stars from becoming carbon star before leaving the AGB. It also sets an upper limit to the distance modulus of the Large Magellanic Cloud of 18.8 mag.

The equivalent width of the absorption band decreases with redder (K − L) colour when the dust continuum emission becomes stronger than the photospheric emission. Carbon stars with similar (K − L) appear to have equally strong 3 μm absorption in the MCs and the Milky Way. We discuss the implications for the carbon and nitrogen enrichment of the stellar photosphere of carbon stars.

Accepted by: Astronomy &Astrophysics Main Journal

For preprints, contact jvanloon@astro.uva.nl
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9902284

On the Spatial Distribution of Stellar Populations in the Large Magellanic Cloud

J. Harris and D. Zaritsky^{1}

1 Dept. of Astronomy and Astrophysics, UC Santa Cruz, USA

We measure the angular correlation function of stars in a region of the Large Magellanic Cloud (LMC) that spans 2.0^{∘} × 1.5^{∘}. We find that the correlation functions of stellar populations are represented well by exponential functions of the angular separation for separations between 2 and 40 arcmin (corresponding to ∼ 30 pc and 550 pc for an LMC distance of 50 kpc). The inner boundary is set by the presence of distinct, highly correlated structures, which are the more familiar stellar clusters, and the outer boundary is set by the observed region’s size and the presence of two principal centers of star formation within the region. We also find that the normalization and scale length of the correlation function changes systematically with the mean age of the stellar population. The existence of positive correlation at large separations (∼ 300 pc), even in the youngest population, argues for large-scale hierarchical structure in current star formation. The evolution of the angular correlation toward lower normalizations and longer scale lengths with stellar age argues for the dispersion of stars with time. We show that a simple, stochastic, self-propagating star formation model is qualitatively consistent with this behavior of the correlation function.

Accepted by: The Astronomical Journal For preprints, contact jharris@ucolick.org

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9902302

Star Formation Histories from HST Color-Magnitude Diagrams of Six Fields of the LMC

K.A.G. Olsen^{1}

1 Cerro Tololo Inter-American Observatory, Casilla 603, La Serena, Chile

We present results on the analysis of background field stars found in HST WFPC2 observations of six of the old globular clusters of the Large Magellanic Cloud. Treated as contaminants by the globular cluster analysts, we produce V − I, V color-magnitude diagrams of the field stars and use them to explore the LMC’s star formation history. The photometry approaches V ∼ 26, well below the turnoff of an ancient (∼14 Gyr) LMC population of stars. The field star CMDs are generally characterized by an upper main sequence broadened by stellar evolution, an old red giant branch, a prominent red clump, and an unevolved lower main sequence. The CMDs also contain a few visual differences, the most obvious of which is the smeared appearance of the NGC 1916 field caused by heavy differential reddening. More subtly, the base of the subgiant branch near the old turnoff appears extended in V and the red giant branch appears broad in V − I in four of the fields, but not in the NGC 1754 field. We use a maximum-likelihood technique to fit model CMDs drawn from Bertelli et al. (1994) isochrones to the observed CMDs. We constrain the models by the age-metallicity relation derived from LMC clusters, test four IMF slopes, and fit for the reddening, distance modulus, and star formation rate. We find that we can just resolve structure in SFR(t) with time steps of ∼0.15 in log age, implying a resolution of ∼ 4 × 10^{8} years at an age of 1 Gyr. For a Salpeter IMF, our derived star formation history for the NGC 1754 field is characterized by an enhanced star formation rate over the past 4 Gyr, qualitatively resembling that derived by others for a variety of LMC fields. The remaining four fields, however, appear to have had high levels of star formation activity as long as 5−8 Gyr ago; these fields lie in the LMC Bar while the NGC 1754 field lies in the disk, suggesting that the inner regions of the LMC contain significantly more older stars than the outer regions. Examining the residuals of the models and observations, we find that the old red giant branches of the models provide a poor fit to the observations, which suggests an error in the model isochrones. The effect of the disagreement appears to be to underestimate the contribution of the old population.

Accepted by: the Astronomical Journal For preprints, contact kolsen@noao.edu

Also available from the URL http://www.ctio.noao.edu/∼olsen/
or by anonymous ftp at ftp://ftp.ctio.noao.edu/pub/olsen/

Constraints on Intervening Stellar Populations Toward the Large Magellanic Cloud

Dennis Zaritsky^{1}, Stephen A. Shectman^{2}, Ian Thompson^{2}, Jason Harris^{1}, & D.N.C. Lin^{1}

1 UCSC, Santa Cruz, USA 2 Carnegie Observatories, Pasadena, USA

The suggestion by Zaritsky & Lin (1997; ZL) that a vertical extension of the red clump feature in color-magnitude diagrams (CMDs) of the Large Magellanic Cloud (LMC) is consistent with a significant population of foreground stars to the LMC that could account for the observed microlensing optical depth (Renault et al. 1997; Alcock et al. 1997a) has been challenged by various investigators (cf. Alcock et al. 1997b, Gallart 1998, Bennett 1998, Gould 1998, Beaulieu & Sackett 1998, and Ibata, Lewis, & Beaulieu 1998). We respond by (1) examining each of the challenges presented to determine whether any or all of those arguments invalidate the claims made by ZL and (2) presenting new photometric and spectroscopic data obtained in an attempt to resolve this issue. We systematically discuss why the objections raised so far do not unequivocally refute ZL’s claim. We conclude that although the CMD data do not mandate the existence of a foreground population, they are entirely consistent with a foreground population associated with the LMC that contributes significantly (∼ 50%) to the observed microlensing optical depth. From our new data, we conclude that ≤ 40% of the VRC stars are young, massive red clump stars because (1) synthetic color-magnitude diagrams created using the star formation history derived independently from HST data (Geha et al. 1998) suggest that < 50% of the VRC stars are young, massive red clump stars, (2) the angular distribution of the VRC stars is more uniform than that of the young (age < 1 Gyr) main sequence stars, and (3) the velocity dispersion of the VRC stars in the region of the LMC examined by ZL, 18.4 ± 2.8 km sec^{−1} (95% confidence limits), is inconsistent with the expectation for a young disk population. Each of these arguments is predicated on assumptions and the conclusions are uncertain. Therefore, an exact determination of the contribution to the microlensing optical depth by the various hypothesized foreground populations, and the subsequent conclusions regarding the existence of halo MACHOs, requires a detailed knowledge of many complex astrophysical issues, such as the IMF, star formation history, and post-main sequence stellar evolution.

Accepted by: The Astronomical Journal For preprints, contact dennis@ucolick.org

Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9902098

A New Argument Against An Intervening Stellar Population Toward the Large Magellanic Cloud

Andrew Gould^{1}

1 Ohio State University, Columbus, USA

Zaritsky & Lin have claimed detection of an intervening population of stars toward the Large Magellanic Cloud (LMC) which, they believe, could account for a substantial fraction of the observed microlensing events. I show that the observed time scales of these events imply that if such an intervening population existed and gave rise to a significant fraction of the microlensing events, then it could not be associated with the LMC. Hence, the radial velocity of the putative intervening population should differ from that of the LMC by of order 100 km s^{−1}. The fact that the radial velocities of the two populations are consistent within errors is therefore strong evidence that these intervening stars do not trace a population that is responsible for the microlensing events.

Submitted to: The Astrophysical Journal

For preprints, contact gould@astronomy.ohio-state.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9902374

A Complete Set of Solutions For Caustic-Crossing Binary Microlensing Events

M.D. Albrow ^{1}, J.-P. Beaulieu ^{2}, J.A.R. Caldwell ^{3}, D.L. DePoy ^{4}, M. Dominik ^{2}, B.S. Gaudi ^{4}, A. Gould ^{4}, J. Greenhill ^{5}, K. Hill ^{5}, S. Kane ^{5,6}, R. Martin ^{7},

J. Menzies ^{3}, R.M. Naber ^{2}, R.W. Pogge ^{4} K.R. Pollard ^{1}, P.D. Sackett ^{2}, K.C. Sahu ^{6},

P. Vermaak ^{3}, R. Watson ^{5}, A. Williams ^{7}, (The PLANET Collaboration)

1 Univ. of Canterbury, Dept. of Physics & Astronomy, Private Bag 4800, Christchurch, New Zealand 2 Kapteyn Astronomical Institute, Postbus 800, 9700 AV Groningen, The Netherlands 3 South African Astronomical Observatory, P.O. Box 9, Observatory 7935, South Africa 4 Ohio State University, Department of Astronomy, Columbus, OH 43210, USA 5 Univ. of Tasmania, Physics Dept., G.P.O. 252C, Hobart, Tasmania 7001, Australia 6 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD. 21218, USA 7 Perth Observatory, Walnut Road, Bickley, Perth 6076, Australia

We present a method to analyze binary-lens microlensing light curves with one well-sampled fold caustic crossing. In general, the surface of χ^{2} shows extremely complicated behavior over the 9parameter space that characterizes binary lenses. This makes it difficult to systematically search the space and verify that a given local minimum is a global minimum. We show that for events with well-monitored caustics, the caustic-crossing region can be isolated from the rest of the light curve and easily fit to a 5-parameter function. Four of these caustic-crossing parameters can then be used to constrain the search in the larger 9-parameter space. This allows a systematic search for all solutions and thus identification of all local minima. We illustrate this technique using the PLANET data for MACHO 98-SMC-1, an excellent and publicly available caustic-crossing data set. We show that a very broad range of parameter combinations are compatible with the PLANET data set, demonstrating that observations of binary-lens lightcurves with sampling of only one caustic crossing do not yield unique solutions. The corollary to this is that the time of the second caustic crossing cannot be reliably predicted on the basis of early data including the first caustic crossing alone. We investigate the requirements for determination of a unique solution and find that occasional observations of the first caustic crossing may be sufficient to derive a complete solution.

Submitted to: The Astrophysical Journal

For preprints, contact gould@astronomy.ohio-state.edu
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9903008

Distances, ages and epoch of formation of Globular Clusters

E. Carretta^{1}, R.G. Gratton^{1}, G. Clementini^{2}, F. Fusi Pecci^{2,3}

1 Osservatorio Astronomico di Padova, Italy 2 Osservatorio Astronomico di Bologna, Italy 3 Stazione Astronomica di Cagliari, Italy

We review the results on distances and absolute ages of galactic globular clusters (GCs) obtained after the release of the Hipparcos catalogue. Several methods aimed at the definition of the Population II local distance scale are discussed. While some methods yield distances to the center of the LMC (the bar), other give distances to the system of globular clusters. There is some rough evidence that the two distances may not coincide. Emphasis is given to the discussion of distances and ages of GCs derived using Hipparcos parallaxes of local subdwarfs. Distances and ages for the 9 clusters discussed in Gratton et al 1997 (ApJ 491, 749, Paper I) are re-derived using an enlarged sample of local subdwarfs, which includes about 90% of the metal-poor dwarfs with accurate parallaxes (Δπ/π ≤ 0.12) in the whole Hipparcos catalogue. On average, our revised distance moduli are decreased by 0.04 mag with respect to Paper I. The corresponding age of the globular clusters is t = 11.8 ± 2.6 Gyr. The corresponding (true) LMC distance modulus (μ_{LMC} = 18.55 ± 0.12 mag for the clusters, and μ_{LMC} = 18.66 ± 0.12 mag for the bar) is compared with other existing determinations. We conclude that at present the best estimate for the distance to the bar of the LMC is: μ_{LMC} = 18.54±0.04 suggesting that distances from the Subdwarf fitting method are 1 σ too long. Our best estimate for the age of the globular clusters is thus revised to: Age = 13.2 ± 2.9 Gyr (95% confidence range). The best relation between ZAHB absolute magnitude and metallicity is: M_{V}(ZAHB) = (0.18 ± 0.09)([Fe/H]+1.5)+(0.65±0.11).

Submitted to: Astrophysical Journal

For preprints, contact carretta@pdmida.pd.astro.it, gisella@astbo4.bo.astro.it
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9902086
or via anonymous ftp at boas5.bo.astro.it/pub/gisella/papers/paper2.ps

Predicted HST FOC and broad band colours for young and intermediate Simple Stellar Populations

E. Brocato^{1,2}, V. Castellani^{3}, G. Raimondo ^{1,2}, and M. Romaniello^{4,5,6}

1 Osservatorio Astronomico di Collurania, Via M. Maggini, I-64100 Teramo, Italy 2 Istituto Nazionale di Fisica Nucleare, LNGS, I-67100 L’Aquila, Italy 3 Dipartimento di Fisica dell’Università di Pisa, Piazza Torricelli 2, I-56126 Pisa, Italy 4 Scuola Normale Superiore, Piazza dei Cavalieri 7, I-56126 Pisa, Italy 5 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 6 European Southern Observatory, Karl-Schwarzschildstr. 2, Garching b. München, D–85748, Germany

This paper presents theoretical HST and broad band colours from population synthesis models based on an homogeneous set of stellar evolutionary tracks as computed under canonical (no overshooting) assumptions, covering the range of cluster ages from t = 8 Myr to t = 5 Gyr for three different metallicities (Z = 0.02, 0.006, and 0.001). Statistical fluctuations in the cluster population have been investigated, assessing the predicted fluctuations of the various colours as a function of the cluster integrated absolute magnitude. We show that the red leak in HST UV filters deeply affects the predicted fluxes and colours. However, we find that for F152M−F307M≤ 0.5 and for F170M−F278M≤ 0.5 (which means ages lower than 1 Gyr) the HST UV colours can still be used to infer reliable indications on the age of distant clusters. Moreover, one finds that the age calibration of these colours is scarcely affected by the amount of original helium or by the assumed IMF. On this basis, we present a calibration of the HST UV two-colours (F152M−F307M vs F170M−F278M) in terms of cluster ages for the three above quoted metallicities. We suggest the combined use of HST UV colours and IR colours (V−K in particular) to disentangle the metallicity-age effect in integrated colours of young stellar populations (t ≤ 1Gyr).

Accepted by: Astronomy & Astrophysics Supplement Series

For preprints, contact brocato@astrte.te.astro.it
Also available from the URL http://xxx.lanl.gov/abs/astro-ph/9812323

Starburst99: Synthesis Models for Galaxies with Active Star Formation

Claus Leitherer^{1}, Daniel Schaerer^{2}, Jeffrey D. Goldader^{3} Rosa M. González Delgado ^{4}, Carmelle Robert^{5}, Denis Foo Kune ^{1,6} Duília F. de Mello^{1}, Daniel Devost^{1,5}, Timothy M. Heckman^{7}

1 Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 2 Observatoire Midi-Pyrénées, 14, Av. E. Belin, F-31400 Toulouse, France 3 Univ. of Pennsylvania, Dept. of Physics & Astronomy, Philadelphia, PA 19104-6396, USA 4 Inst. Astrofísica de Andalucía, Apartado 3004, E-18080 Granada, Spain 5 Département de Physique and Observatoire du mont Mégantic, Université Laval, Québec, QC, G1K 7P4, Canada 6 Macalester College, Dept. of Physics & Astronomy, 1600 Grand Ave., St. Paul, MN 55105, USA 7 Physics and Astronomy Department, Johns Hopkins University, Homewood Campus, Baltimore, MD 21218, USA

Starburst99 is a comprehensive set of model predictions for spectrophotometric and related properties of galaxies with active star formation. The models are an improved and extended version of the data set previously published by Leitherer & Heckman (1995). We have upgraded our code by implementing the latest set of stellar evolution models of the Geneva group and the model atmosphere grid compiled by Lejeune et al. (1997). Several predictions which were not included in the previous publication are shown here for the first time. The models are presented in a homogeneous way for five metallicities between Z = 0.040 and 0.001 and three choices of the initial mass function. The age coverage is 10^{6} to 10^{9} yr. We also show the spectral energy distributions which are used to compute colors and other quantities. The full data set is available for retrieval at http://www.stsci.edu/science/starburst99/. This website allows users to run specific models with non-standard parameters as well. We also make the source code available to the community.

Accepted by: Astrophysical Journal Supplement Series

For preprints, contact Claus Leitherer (leitherer@stsci.edu)
Also available from the URL http://www.stsci.edu/science/starburst99

Abstracts of Non-Refereed Papers

What Could the Machos Be?

HongSheng Zhao^{1}

1 Sterrewacht Leiden, Niels-Bohrweg 2, 2333 CA Leiden, The Netherlands

If the Universe has a significant baryonic dark component in the form of compact objects in galaxy halos (machos), then there is a minute chance (about 10^{−7}) that one of the Galactic machos passes sufficiently close to our line of sight to a star out of some 10^{7} monitored stars in the Magellanic Clouds (MCs) that it brightens by more than 0.3 magnitude due to gravitational focusing. After a brief discussion of the current controversy over the interpretation of the observed events, i.e., whether the lensing is caused by halo white dwarfs or machos in general or by stars in various observed or hypothesized structures of the Clouds and the Galaxy, I propose a few observations to put ideas of the pro-macho camp and the pro-star camp to test. In particular, I propose a radial velocity survey towards the MCs.

To appear in: AIP Proc. “Particle Physics and the Early Universe” (COSMO-98, Asilomar, California, November 1998), ed. David Caldwell

For preprints, contact hsz@strw.leidenuniv.nl
Also available from the URL http://babbage.sissa.it/abs/astro-ph/9902179

Job Opportunity

Postdoctoral Fellowship at the University of Bonn

The Graduiertenkolleg “The Magellanic Clouds and Other Dwarf Galaxies”, jointly run at the Universities of Bonn and Bochum, funded by the Deutsche Forschungsgemeinschaft, invites applications for a postdoctoral fellowship. The Fellow is meant to work at Bonn University.

The successful applicant will find her/himself in a pleasant scientific environment, closely collaborating with graduate students on related topics. She/he should be interested in galaxy research, in particular in the investigation of dwarf galaxies. Members of the Graduiertenkolleg currently pursue both theoretical and observational studies in the field of galaxy research. They have regular access to major radio and optical telescopes around the world, and also utilize satellite observatories.

The major goal of this graduate or research school is to furnish comprehensive scientific experience and education of graduate students in astrophysics, while working in a team. The scientific focus is on the star formation history and evolution, on the interstellar medium, and on the structure and kinematics of the Magellanic Clouds and of more distant dwarf galaxies.

For further information, please, see also our grad school home page:

                     http://www.astro.uni-bonn.de/˜ webgk

The fellowship will be available as of 1st of July, 1999. The appointment is for one year with extension up to a total of 2 years. An appointment for a shorter, work-visit like period is also possible.

The monthly payment, set by the rules of the DFG, is DM 2490.- and 2690.- (depending on age), plus DM 200.- for expenditures, plus a DM 400.- allowance for married persons.

Applications should be sent (by 15 April 1999) to the Spokesman of the Graduiertenkolleg:

Prof. Dr. Klaas S. de Boer Sternwarte der Universität Bonn Auf dem Hügel 71 D-53121 Bonn, Germany.

deboer@astro.uni-bonn.de